### v3

# The Prague Relational Learning Repository

Abstract
The aim of the Prague Relational Learning Repository is to support machine learning research with multi-relational data. The repository currently contains 148 SQL databases hosted on a public MySQL server located at relational.fel.cvut.cz. The server is provided by CTU, the Czech technical university, to support the relational machine learning community. A searchable meta-database provides metadata (e.g., the number of tables in the database, the number of rows and columns in the tables, the number of self-relationships).
1 Goals
Many organizations maintain their data in relational databases, which support complex structured data. Extending machine learning from traditional single-table methods to multi-relational data is an important direction for practical applications. The statistical and algorithmic challenges that arise from multi-relational data have been addressed in a number of research communities, such as Statistical-Relational Learning, Multi-Relational Data Mining, and Inductive Logic Programming. Experience with the UCI Machine Learning Repository [^1] http://archive.ics.uci.edu/ml/ has shown that a shared repository of benchmark datasets facilitates research progress [1]. The UCI Machine Learning Repository contains mainly datasets stored in a single data table. Our goal is to provide a similar service for the relational learning community for relational datasets that contain multiple interrelated tables.
2 Design
The repository is maintained in a public MySQL server hosted by Czech Technical University in Prague. Each dataset is stored as a MySQL database on the server. Different formats have been introduced for storing multi-relational data. The advantages of using the SQL (SQL stands for “Structured Query Language”) format include the following.
  • The SQL format is a based on a standard widely used in industry. Using SQL databases in machine learning facilitates cross-community knowledge transfer and collaborations between machine learning and database researchers.
  • Because SQL is a common standard, many programming environments support accessing and processing SQL data. This includes machine learning and statistical platforms such as Clowdflows [13], RapidMiner, and Weka. All general application languages provide SQL database connectivity, including R, Python, Java, and C++.
  • The data description facilities of SQL provide a standard for defining metadata about the structure of the dataset. For example, information about the entities linked by a relationship is specified using primary and foreign keys. This metadata is recorded in the system catalog, and can be queried by machine learning applications.
To facilitate using tools developed for other relational data formats, we have provided scripts for converting MySQL data to other common data formats used in relational learning. [^2] https://www2.cs.sfu.ca/~oschulte/jbn/DataConversion/MLN.html This includes the Wisconsin Logic Learning format (WILL) and the .db format used in the Alchemy system. The ClowdFlows system also provides data format conversion, for example from MySQL to the Aleph Inductive Logic Programming Format. [^3] http://www.cs.ox.ac.uk/activities/machlearn/Aleph/aleph.html
3 Content
The repository currently contains 148 databases. This includes common benchmark datasets used in relational learning, like eastbound/westbound train dataset [15] or biodegradability dataset [3]. Some of the databases are derived from the same base data in different ways (e.g. the repositor contains different version of the IMDb dataset). We have aimed at providing a diversity of databases, for instance in terms of the number of records and in terms of the complexity of the relational schema. Hence, also synthetic datasets from different database vendors are included, as they are designed to show off capabilities of their database software. An example of such a synthetic dataset is Adventure Works, which is interesting not only because of its complexity, but also because of:
  • it uses both, simple and composite keys;
  • it contains a diverse set of data types, including datetime, blob (images) and geometry;
  • it contains missing values.
4 Access and Contributions
Read-only access can be obtained via a database connection with the following parameters.
Hostname
relational.fel.cvut.cz
Port
3306
Username
guest
Password
ctu-relational
To contribute a database, please contact the repository administrators; a web contact form is available https://relational.fel.cvut.cz/contact. One possibility is to provide us with a MySQL dump of your database. Another option is to provide us with read access to your database on your server, so we can migrate the database to the public server.
5 The Meta-Database
Table 1 shows selected metadata from the meta-database. The meaning of the columns is as follows.
#Relations
The number of tables in the database.
#Instances
Count of rows in the target table.
Size
Size in MB including indexes.
Type
The dataset is either a measurement or synthetically generated.
Domain
The original domain.
Task
Classification or regression.
Database #Relations #Instances Size Type Domain Task
Accidents [18] 4 495760 210.0 Real Goverment Class
AdventureWorks 71 30669 234.6 Synth Retail Regr
AustralianFootball 4 3036 38.0 Real Sport Class
Biodegradability [3] 5 328 3.2 Real Medicine Regr
Carcinogenesis [24] 6 329 26.3 Real Medicine Class
CCS 6 1000 658.4 Real Finance Regr
ClassicModels 8 273 0.5 Synth Retail Regr
Countries 4 247 8.6 Real Geography Regr
Credit 9 10084 443.6 Synth Retail Class
CS 8 100 0.3 Synth Finance Class
Dunur [11] 20 276 0.8 Real Kinship Class
Elti [11] 14 1081 0.7 Real Kinship Class
Employee 7 2838426 344.6 Synth Retail Regr
Financial [2] 8 682 94.1 Real Finance Class
FTP 2 29555 7.5 Synth Retail Class
Genes [6] 3 862 1.9 Real Medicine Class
Hepatitis 7 500 2.2 Real Medicine Class
Hockey [23] 23 7759 15.5 Real Sport Class
IMDb [23] 7 794625 614.6 Real Entertainment Class
MovieLens [22] 7 6039 151.9 Real Entertainment Class
Lahman 25 23111 84.0 Real Sport Regr
LegalActs 5 564268 238.2 Real Gouverment Class
Mesh [9] 32 223 1.1 Real Industry Regr
Mondial [14] 33 454 3.3 Real Geography Class
MooneyFamily [20] 72 92 3.3 Synth Kinship Class
Mutagenesis [8] 3 188 0.9 Real Medicine Class
Nations 3 14 2.1 Real Geography Class
NBA [23] 4 30 0.3 Real Sport Class
NCAA 10 268 40.6 Real Sport Class
Northwind 29 830 1.1 Synth Retail Regr
Pima 14 768 0.8 Real Medicine Class
PremiereLeague [19] 4 363 11.3 Real Sport Class
PTE [25] 41 299 7.3 Real Medicine Class
Pubs 11 18 0.4 Synth Retail Regr
Sakila 16 15991 6.6 Synth Retail Regr
SalesDB 4 6148886 539.3 Synth Retail Regr
SameGen 7 1081 0.3 Real Kinship Class
Stats 8 38357 621.4 Real Education Regr
StudentLoan [17] 13 1000 0.9 Real Education Class
PTC [10] 4 343 7.8 Real Medicine Class
Thrombosis [7] 3 806 1.9 Real Medicine Class
TPCC [12] 9 28433 174.1 Synth Retail Class
TPCDS [16] 24 99550 4587.5 Synth Retail Class
TPCH [4] 8 148255 1925.1 Synth Retail Regr
Trains [15] 2 20 0.1 Synth Logistic Class
University 5 38 0.3 Synth Education Class
UW-CSE [21] 4 278 0.2 Real Education Class
VOC 8 8215 2.7 Real Logistic Class
World 3 239 0.8 Real Geography Class
Table 1: List of databases in the repository
The name of the meta-database schema is meta. This schema contains a number of tables with information about the databases, as well as the performance of different learning algorithms on the databases. The name of the table that contains information about the databases is meta.information. Some of this metadata is automatically exported in HTML format for display on the webpage relational-data.org. In the following, we list the names of the main column and their meaning. When we refer to “all columns” or “all rows”, we mean all columns/rows of all tables in a database. The metadata contain the following main groups of information: basic database statistics, information about columns or fields, foreign key structure, classification information.
5.1 Basic Database Statistics
Various basic properties, such as record count and missing values.
row_count
The total number of rows, or records.
row_max
The maximum number of rows, or records, in a single table.
column_count
The total number of columns, or fields.
download_url
A URL containing further information about the dataset, such as provenance.
null_count
The number of table entries with null values; typically this is the number of table entries with missing values.
5.2 Column Information
These columns contain metainformation about the types of columns/fields/attributes in the database tables. The list is mutually exclusive and collectively exhaustive as it holds: $column\_count=geo\_count+date\_count+lob\_count+string\_count+numeric\_count$.
geo_count
The number of columns that represent spatial attributes. (These are called “geographic” features in MySQL.)
date_count
The number of columns that represent temporal attributes (date, time, or year).
lob_count
The number of columns that store large objects (e.g., images).
string_count
The number of columns that store string values. This typically includes discrete attributes.
numeric_count
The number of numeric columns.
5.3 Foreign Key Structure
A foreign key points from one table to another. Chen et al. propose visualizing the foreign key relationships in a semantic relationship graph [5]: The graph contains a directed edge from table $T$ to table $T^{\prime}$ if table $T$ references $T^{\prime}$ in a foreign key constraint. These columns represent information about the structure of the semantic relationship graph.
primary_key_count
The number of primary keys.
composite_key_count
The number of primary keys that comprise more than one column.
foreign_key_count
The number of foreign keys.
self_referencing_table_count
The number of tables such that the table contains a foreign key pointer to one of its own columns. This occurs for example when a relational schema represents a class hierarchy or taxonomy.
has_loop
Whether there exists a loop of foreign key pointers over several tables. An example of a loop is when between a person table and a university table exists two foreign keys - the first foreign key signifies that a person is studying at a university, while the second foreign key signifies that the person is teaching at the university.
5.4 Classification
Many of the databases in the repository have been used to study classification in relational data. There is often a standard class label for such studies; we refer to this as the target attribute. These columns contain information relevant to the target attribute where it exists.
target_column
The target attribute most often used in relational classification studies.
target_table
The table that contains the target column.
target_id
The primary key field of the target table.
instance_count
The number of rows in the target table.
class_count
The number of class labels.
majority_class_ratio
The proportion of the majority class label on instance count.
6 Conclusions
In this paper, we presented the Prague Relational Learning Repository (PRLR), an easily accessible collection of datasets for relational learning. The PRLR was designed with supervised learning in mind. To this end, the PRLR contains 148 ready to download datasets. One of the important features of the PRLR is that it provides meta-data about the datasets. The PRLR meta-data can be accessed at https://relational.fel.cvut.cz.
Acknowledgment
We are grateful to CTU Prague for sponsoring and maintaining the database server. Our industry partner getML ensured service continuity by hosting the database server during a transition phase (www.getml.com). We would like to thank all of the donors who contributed data to the repository, and the author of the web page, Václav Ostrožlík. This work was supported by the Grant Agency of the Czech Technical University in Prague, grant No. SGS15/117/OHK3/1T/18.